OpenStudy (anonymous):

Evaluate the function at the specified values of the independent variable and simplify

OpenStudy (anonymous):

f(x)= sqrt(x+8)+2 (a) f(-8) (b) f(1) (c) f(x-8)

OpenStudy (alexwee123):

looks like a simple plug in for A) sub in -8 for x and just simplify

OpenStudy (anonymous):

What about b and C? Also is this right f(x) = sqrt(x+8) + 2 f(-8) = sqrt(-8+8) + 2 f(-8) = sqrt(0) + 2 f(-8) = 2

OpenStudy (anonymous):

thats correct b)\[f(1)=\sqrt{1+8}+2=\sqrt{9}+2=3+2=5\] c)\[f(x-8)=\sqrt{x-8+8}+2=\sqrt{x}+2=x+2\]